ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 136, 2003

Donin, J. and A. Mudrov Reflection equation, twist, and equivariant quantization

11

Eisworth, T., P. Nyikos and S. Shelah Gently killing S-spaces  189

Emel’yanov, E. Y. Invariant densities and mean ergodicity of Markov operators

373

Felix, Y. and P.-E. Parent Density and unique decomposition theorems for the lattice of cellular classes

317

Fonf, V. P. and J. Lindenstrauss Boundaries and generation of convex sets

157

Friedland, S. and A. S. Libgober Generalizations of the odd degree theorem and applications

353

Furman, A. On minimal, strongly proximal actions of locally compact groups

173

Guo, X. Y. and K. P. Shum Permutability of minimal subgroups and p-nilpotency of finite groups

145

Koskela, P. and K. Rajala Mappings of finite distortion: Removable singularities

269

Letzter, E. S. Noetherianity of the space of irreducible representations

307

Libgober, A. S. See S. Friedland

Lindenstrauss, J. See V. P. Fonf

Loya, P. Asymptotic properties of the heat kernel on conic manifolds 285

Mudrov, A. See J. Donin

Nyikos, P. See T. Eisworth

Pang, X. and L. Zalcman Normal families of meromorphic functions with multiple zeros and poles

1

Parent, P.-E. See Y. Felix

Rajala, K. See P. Koskela

Rugeihyamu, S. E. On the Bieberbach and Koebe constants for sector domains and sector disks

117

Shelah, S. Not collapsing cardinals ≤ κ in (< κ)-support iterations  29

Shelah, S. See T. Eisworth Shum, K. P. See X. Y. Guo

Tsunogai, H. The stable derivation algebras for higher genera  221

Wen, Z.-Y. and L.-F. Xi Relations among Whitney sets, self-similar arcs and quasi-arcs

251

Xi, L.-F. See Z.-Y. Wen Zalcman, L. See X. Pang Zhao, Y. Zeros of Abelian integrals for the reversible codimension four

quadratic centers Q_{3}^{R}∩Q_{4} 125